# The R&D Productivity Puzzle: Innovation Networks with Heterogeneous Firms

### v2

@@ Abstract
We introduce heterogeneous R&D productivities into an endogenous R&D network formation model, generalizing the framework of Goyal and Moraga-González (2001). Heterogeneous productivities endogenously create asymmetric gains from collaboration: less productive firms benefit disproportionately from links, while more productive firms exert greater R&D effort and incur higher costs. When productivity gaps are sufficiently large, more productive firms experience lower profits from collaborating with less productive partners. As a result, the complete network—stable under homogeneity—becomes unstable, and the positive assortative (PA) network, in which firms cluster by R&D productivity, emerges as pairwise stable. Using simulations, we show that the clustered structure delivers higher welfare than the complete network; nevertheless, welfare under this formation follows an inverted U-shape as the fraction of high-productivity firms increases, reflecting crowding-out effects at high fractions. Altogether, we uncover an R&D productivity puzzle: economies with higher average R&D productivity may exhibit lower welfare through (i) the formation of alternative stable networks, or (ii) a crowding-out effect of high-productivity firms. Our findings highlight that productivity is a structural force reshaping the organization of innovation. Productivity-enhancing policies must therefore account for their effects on endogenous R&D alliances and effort, as they may reverse the intended welfare gains. JEL Classification: D21, D85, O30.
@@ 1 Introduction
In this paper, we study R&D alliances and their implications for firms’ profits and aggregate welfare in an R&D network formation model with heterogeneous R&D productivities. A growing theoretical literature, including Goyal and Moraga-González (2001); Goyal and Joshi (2003); König et al. (2019); Hsieh et al. (2025), studies R&D networks where firms improve their R&D efficiencies by forming collaborations while still competing with each other in the product market. However, these models typically assume that firms are homogeneous in their ability to transform R&D effort into cost-reducing innovations. While analytically convenient, this assumption is at odds with theoretical models and empirical evidence showing that capabilities vary substantially across firms (Magerman et al., 2016; Chen et al., 2023; Milán, 2019; Acemoglu and Tahbaz-Salehi, 2025). [^1] Further empirical studies show that the return on R&D is driven by unobserved heterogeneity in innovation efficiency (Lentz and Mortensen, 2008) and organizational competence (Henderson and Cockburn, 1994), rather than spending intensity alone. Indeed, the output elasticity of R&D is highly firm-specific (Knott, 2008) and depends on the interplay between a firm’s size and innovation strategy (Akcigit and Kerr, 2018). We show that once this assumption is relaxed and firms are heterogeneous, a productivity puzzle is introduced: economies with larger overall R&D productivity can exhibit lower welfare.
@@ 2.1.2 Equilibrium R&D Efforts
$\left.\frac{\partial{\pi}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)}{\partial e_{i}}\right|_{{e}^{*}_{i}}=\quad\frac{2{q}^{*}_{i}(\mathbf{e},\mathbf{G})}{n+1}\left(-n\left.\frac{\partial c_{i}(\mathbf{e},\mathbf{G})}{\partial e_{i}}\right|_{{e_{i}}^{*}}+\sum_{j\neq i}\left.\frac{\partial c_{j}(\mathbf{e},\mathbf{G})}{\partial e_{i}}\right|_{{e_{i}}^{*}}\right)-2\phi{e}^{*}_{i}=0,$
@@
$\phi>\underline{\phi}\coloneqq\frac{n^{3}-n^{2}+n}{(n+1)^{2}}.$ (8)
Proposition 1 establishes that a unique and interior equilibrium exists for any network $\mathbf{G}\in\mathcal{G}^{n}$, provided the cost coefficient $\phi$ exceeds a threshold $\underline{\phi}$ that grows linearly with $n$. The proof of Proposition 1 relies on the Levy-Desplanques theorem, itself a consequence of Gershgorin’s disc theorem (Gershgorin, 1931). Specifically, we provide a bound for $\phi$ that guarantees the matrix $\mathbf{A}(\mathbf{G})$ is strictly diagonally dominant regardless of the network $\mathbf{G}$, which implies that it is non-singular and hence the FOC system admits a unique solution.
While existence and uniqueness of an equilibrium effort profile are guaranteed by Proposition 1, equilibrium effort levels cannot be generally expressed in closed form, as they involve an intractable inverse of the non-symmetric matrix $\mathbf{A}(\mathbf{G})$. Nevertheless, we can derive equilibrium profits using equilibrium efforts as:
@@
This can be compared with the result in Hsieh et al. (2025), showing ${e}^{*}_{i}=\theta{q}^{*}_{i}/(2\phi)$ (with the relevant changes in notation). The differences in equilibrium effort and profit levels in these two setups arise from, first, differences in sources of heterogeneity (in baseline costs in their model vs. R&D productivities in ours) and, second, the solution methods implemented for equilibrium construction (simultaneous in their model vs. sequential action in ours). Equation (11) implies that a firm’s equilibrium size is positively related to its equilibrium R&D effort, consistent with empirical findings in Cohen and Klepper (1996). Additionally, we see from the expression that the optimal output level is negatively related to its productivity $\theta_{i}$ and its sparsity coefficient $\eta_{i}(\mathbf{G})$. Equations (10) and (11) imply that for comparing one high- to one low-productive firm with $\theta_{i}>\theta_{j}$:
@@ 2.1.3 R&D Network Formation
Pairwise stability requires that no firm has an incentive to sever an existing link, and that no two firms both benefit (with at least one strictly gaining) from forming a new one (Jackson and Wolinsky, 1996). In our case, the formal statement translates to the following: a network $\mathbf{G}$ is pairwise stable if and only if for all firms $i,j\in\mathcal{N}$,
@@
  • Potential Alliances: If ${G}_{ij}=0$, then at least one of the firms would incur a loss in profit by forming the alliance, i.e., if $\pi_{i}(\mathbf{G}_{+ij})>\pi_{i}(\mathbf{G})$, then $\pi_{j}(\mathbf{G}_{+ij})<\pi_{j}(\mathbf{G})$.
@@ 2.2 Simulation preview: Link sustainability for random networks
Note: Networks $\mathbf{G}$ with $n=20$ firms generated as random binary adjacency matrices according to an Erdos-Renyi scheme with connectivity parameter $\ell\in\{0,0.25,0.75,1\}$, represented by different colors. Productivity distributions $(\theta_{1},\ldots,\theta_{n})$ drawn as random samples from a Beta(0.5,0.5)(0.5,0.5) distribution (left); Beta(1,1)(1,1) (middle); and Beta(2,2)(2,2) (right), with corresponding density plotted in the panels above. The percentage change in profit of firm $i$ (before and after adding a link to firm $j$) is drawn using a solid line, and for firm $j$ using a dashed line. Productivity of the higher-productive firm $i$ is chosen from $\theta_{i}\in\{0.25,0.5,0.75\}$. Percentage change in profits for firm $i$ crosses 0 at a threshold ${\theta}^{*}_{ij}\in(0,1)$ for each productivity distribution and connectivity level cases shown. Figures are limited to the range $0<\theta_{j}<\theta_{i}$ of each configuration to highlight the crossing at 0 for the higher-productive firm $i$.
@@ Appendix A Proofs
We use the Levy-Desplanques theorem, which is a consequence of Gershgorin’s disc theorem (Gershgorin, 1931): an $n\times n$ matrix $A$ is nonsingular if every column is strictly diagonally dominant, i.e. $|A_{ii}|>\sum_{j\neq i}|A_{ij}|$ for all $i=1,\ldots,n$. We will show that a bound on $\phi$ guarantees this property for $\mathbf{A}(\mathbf{G})$ as defined in (7). For any firm $i\in\mathcal{N}$, we have
@@
$\sum_{j\neq i}|A_{ij}(\mathbf{G})|\leq d_{i}(\mathbf{G})\cdot(n-1)+[n-1-d_{i}(\mathbf{G})](n-1)=(n-1)^{2},\quad\text{for each }i\in\mathcal{N}.$
@@
$\min_{i\in\mathcal{N}}|A_{ii}(\mathbf{G})|=\min_{i\in\mathcal{N}}\left|\frac{(n+1)^{2}\phi}{\theta_{i}[n-d_{i}(\mathbf{G})]}-\theta_{i}[n-d_{i}(\mathbf{G})]\right|>(n-1)^{2}.$
Using $\theta_{i}\leq 1$ and $n-d_{i}(\mathbf{G})\leq n$, we see that any $\phi\geq 1$ ensures $A_{ii}(\mathbf{G})>0$. Under these bounds, $\min_{i\in\mathcal{N}}|A_{ii}(\mathbf{G})|\geq[(n+1)^{2}\phi-n^{2}]/n$, and solving for $\phi$ yields the sufficient inequality
$\phi>\frac{n(n-1)^{2}+n^{2}}{(n+1)^{2}}=\frac{n^{3}-n^{2}+n}{(n+1)^{2}}.$
@@
${e}^{*}_{i}(\mathbf{G}^{C})$ $=\frac{\theta_{i}}{(n+1)^{2}\phi-\theta_{i}^{2}}\left[\alpha-\bar{c}+\sum_{j\neq i}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})\right],$
${e}^{*}_{i}(\mathbf{G}^{C})$ $=\frac{\theta_{i}}{(n+1)^{2}\phi}\left[\alpha-\bar{c}+\sum_{j=1}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})\right],$
@@ Appendix B Additional Simulation Results
Note: This figure isolates the effect of changing the productivity composition while holding the network fixed. We consider $n=10$ firms partitioned into two fully connected components (two cliques of size 55), mirroring the connectivity pattern of the Positive Assortative (PA) network at $\rho=\tfrac{1}{2}$. Starting from $\rho=0$ (all firms have low productivity $\theta$), we increase $\rho$ in increments of $0.1$ by upgrading one firm at a time from $\theta$ to 11 (high productivity). At each step, we record the change in profit for the transitioning firm (the firm whose productivity is upgraded in that step), holding all other firms’ productivities and the network links fixed. The x-axis reports $\rho$ (the share of high-productivity firms), and the y-axis reports the corresponding profit change of the transitioning firm. Curves are shown for three values of the low productivity parameter, $\theta\in\{0.1,0.5,0.9\}$. Two patterns emerge. First, the profit gain from upgrading is larger when $\theta$ is smaller, consistent with a larger productivity jump from $\theta$ to 11. Second, the profit gain is generally increasing in $\rho$, except around $\rho=\tfrac{1}{2}$, where the fixed network aligns with exact type clustering (all high types concentrated within one clique and all low types within the other); upgrading a low-type firm within the low-type clique yields a smaller marginal benefit than upgrading the remaining low-type firm in the high-type clique. In all cases the profit change remains positive. Finally, note that this fixed-network exercise is not, in general, a stability analysis: the imposed two-clique network need not be pairwise stable for the corresponding productivity composition, except at $\rho=\tfrac{1}{2}$ (and sufficiently low $\theta$).
@@
Note: Each row corresponds to a different number of firms, ranging from $n=5$ at the top to $n=200$ at the bottom. Each column corresponds to a different proportion of high-productivity firms ($\rho$), increasing from $10\%$ (left) to $90\%$ (right). Orange regions indicate parameter combinations where the PA network is stable; blue regions indicate where the FC network is stable.

### v3

@@ Abstract
We introduce heterogeneous R&D productivities into an endogenous R&D network formation model, generalizing the framework of Goyal and Moraga-González (2001). Heterogeneous productivities endogenously create asymmetric gains from collaboration: less productive firms benefit disproportionately from links, while more productive firms exert greater R&D effort and incur higher costs. When productivity gaps are sufficiently large, more productive firms experience lower profits from collaborating with less productive partners. As a result, the complete network—stable under homogeneity—becomes unstable, and the positive assortative (PA) network, in which firms cluster by R&D productivity, emerges as pairwise stable. Using simulations, we show that the clustered structure delivers higher welfare than the complete network; nevertheless, welfare under this formation follows an inverted U-shape as the fraction of high-productivity firms increases, reflecting crowding-out effects at high fractions. Altogether, we uncover an R&D productivity puzzle: economies with higher average R&D productivity may exhibit lower welfare through (i) the formation of alternative stable networks, or (ii) a crowding-out effect of high-productivity firms. Our findings show that productivity gaps shape the organization of innovation by altering equilibrium R&D alliances and effort. Productivity-enhancing policies must therefore account for these endogenous responses, as they may reverse intended welfare gains. JEL Classification: D21, D85, O30.
@@ 1 Introduction
In this paper, we study R&D alliances and their implications for firms’ profits and aggregate welfare in an R&D network formation model with heterogeneous R&D productivities. Following the canonical work by Goyal and Moraga-González (2001), a growing theoretical literature, including Goyal and Joshi (2003); König et al. (2019); Hsieh et al. (2025), studies R&D networks in which firms raise the effectiveness of R&D through collaboration while continuing to compete in product markets. A common benchmark in this literature abstracts from cross-firm differences in the ability to translate R&D effort into cost-reducing innovations. We extend this framework by allowing firms’ innovation efficiencies to differ, consistent with theory and evidence that capabilities vary substantially across firms (Magerman et al., 2016; Chen et al., 2023; Milán, 2019; Acemoglu and Tahbaz-Salehi, 2025). [^1] Further empirical studies show that the return on R&D is driven by unobserved heterogeneity in innovation efficiency (Lentz and Mortensen, 2008) and organizational competence (Henderson and Cockburn, 1994), rather than spending intensity alone. Indeed, the output elasticity of R&D is highly firm-specific (Knott, 2008) and depends on the interplay between a firm’s size and innovation strategy (Akcigit and Kerr, 2018). We show that introducing heterogeneity generates a “productivity puzzle”: economies with larger overall R&D productivity can exhibit lower welfare.
@@ 2.1.2 Equilibrium R&D Efforts
$\left.\frac{\partial{\pi}^{*}_{i}\left(\mathbf{e},\mathbf{G}\right)}{\partial e_{i}}\right|_{{e}^{*}_{i}}=\quad\left(\left.\frac{2{q}^{*}_{i}(\mathbf{e},\mathbf{G})}{n+1}\right|_{{e}^{*}_{i}}\right)\cdot\left(-n\left.\frac{\partial c_{i}(\mathbf{e},\mathbf{G})}{\partial e_{i}}\right|_{{e_{i}}^{*}}+\sum_{j\neq i}\left.\frac{\partial c_{j}(\mathbf{e},\mathbf{G})}{\partial e_{i}}\right|_{{e_{i}}^{*}}\right)-2\phi{e}^{*}_{i}=0,$
@@
$\phi>\frac{n[2(n-1)^{2}+n]}{(n+1)^{2}}\eqqcolon\underline{\phi}.$ (8)
Proposition 1 establishes that a unique and interior equilibrium exists for any network $\mathbf{G}\in\mathcal{G}^{n}$, provided the cost coefficient $\phi$ exceeds a threshold $\underline{\phi}>2(n-1)$. The proof of Proposition 1 relies on both the Levy-Desplanques theorem (itself a consequence of Gershgorin’s disc theorem, Gershgorin, 1931) and matrix power series (e.g., Chapter 5 of Horn and Johnson, 2012). Specifically, we provide a bound for $\phi$ that guarantees the matrix $\mathbf{A}(\mathbf{G})$ is strictly diagonally dominant (which guarantees its non-singularity) and its inverse has positive row sums regardless of the network $\mathbf{G}$. Both facts together imply the FOC system (7) admits a unique solution with strictly positive R&D efforts.
While existence and positivity of a unique equilibrium effort profile are guaranteed by Proposition 1, equilibrium effort levels cannot be generally expressed in closed form, as they involve an intractable inverse of the non-symmetric matrix $\mathbf{A}(\mathbf{G})$. Nevertheless, we can derive equilibrium profits using equilibrium efforts as:
@@
This can be compared with the result in Hsieh et al. (2025), showing ${e}^{*}_{i}=\theta{q}^{*}_{i}/(2\phi)$ (with the relevant changes in notation). The differences in equilibrium effort and profit levels in these two setups arise from, first, differences in sources of heterogeneity (in baseline costs in their model vs. R&D productivities in ours) and, second, the solution methods implemented for equilibrium construction (simultaneous in their model vs. sequential action in ours). Equation (11) implies that a firm’s equilibrium size is positively related to its equilibrium R&D effort, consistent with empirical findings in Cohen and Klepper (1996). Additionally, we see from the expression that the optimal output level is negatively related to its productivity $\theta_{i}$ and its sparsity coefficient $\eta_{i}(\mathbf{G})$.
Equations (10) and (11) imply that for comparing one high- to one low-productive firm with $\theta_{i}>\theta_{j}$:
@@ 2.1.3 R&D Network Formation
Pairwise stability requires that no firm has an incentive to sever an existing link, and that no two firms both benefit (with at least one strictly gaining) from forming a new one (Jackson and Wolinsky, 1996). In our case, the formal statement translates to the following: a network $\mathbf{G}$ is pairwise stable if and only if for all pairs of firms $(i,j)$ where $i,j\neq i\in\mathcal{N}$:
@@
  • Potential Alliances: If ${G}_{ij}=0$, then at least one of the firms would incur a loss in profit by forming the alliance, i.e., $\pi_{i}(\mathbf{G}_{+ij})<\pi_{i}(\mathbf{G})$ or $\pi_{j}(\mathbf{G}_{+ij})<\pi_{j}(\mathbf{G})$.
@@ 2.2 Simulation preview: Link sustainability for random networks
Note: Networks $\mathbf{G}$ with $n=20$ firms generated as random binary adjacency matrices according to an Erdos-Renyi scheme with connectivity parameter $\ell\in\{0,0.25,0.5,0.75,1\}$, represented by different colors. Productivity distributions $(\theta_{1},\ldots,\theta_{n})$ drawn as random samples from a Beta(0.5,0.5)(0.5,0.5) distribution (left); Beta(1,1)(1,1) (middle); and Beta(2,2)(2,2) (right), with corresponding density plotted in the panels above. We set $\phi=\underline{\phi}$ for all networks. The percentage change in profit of firm $i$ (before and after adding a link to firm $j$) is drawn using a solid line, and for firm $j$ using a dashed line. Productivity of the higher-productive firm $i$ is chosen from $\theta_{i}\in\{0.25,0.5,0.75\}$. Percentage change in profits for firm $i$ crosses 0 at a threshold ${\theta}^{*}_{ij}\in(0,1)$ for each productivity distribution and connectivity level cases shown. Figures are limited to the range $0<\theta_{j}<\theta_{i}$ of each configuration to highlight the crossing at 0 for the higher-productive firm $i$.
@@ Appendix A Proofs
We first show that a bound on $\phi$ guarantees $\mathbf{A}(\mathbf{G})$ is non-singular, which establishes uniqueness of the effort profile ${\mathbf{e}}^{*}(\mathbf{G})$. We use the Levy-Desplanques theorem, which is a consequence of Gershgorin’s disc theorem (Gershgorin, 1931): an $n\times n$ matrix $\mathbf{A}$ is nonsingular if every column is strictly diagonally dominant, i.e. $|A_{ii}|>\sum_{j\neq i}|A_{ij}|$ for all $i=1,\ldots,n$. For any firm $i\in\mathcal{N}$, we have
@@
$s^{*}\coloneqq\max_{i\in\mathcal{N}}\sum_{j\neq i}|A_{ij}(\mathbf{G})|\leq\max_{i\in\mathcal{N}}\{d_{i}(\mathbf{G})\cdot(n-1)+[n-1-d_{i}(\mathbf{G})](n-1)\}\leq(n-1)^{2}.$
@@
$d^{*}\coloneqq\min_{i\in\mathcal{N}}|A_{ii}(\mathbf{G})|=\min_{i\in\mathcal{N}}\left|\frac{(n+1)^{2}\phi}{\theta_{i}[n-d_{i}(\mathbf{G})]}-\theta_{i}[n-d_{i}(\mathbf{G})]\right|>(n-1)^{2}\geq s^{*}.$
Using $\theta_{i}\leq 1$ and $n-d_{i}(\mathbf{G})\leq n$, we see that any $\phi\geq 1$ ensures $A_{ii}(\mathbf{G})>0$. Under these bounds, $d^{*}\geq[(n+1)^{2}\phi-n^{2}]/n$, and solving for $\phi$ yields the sufficient inequality
$\phi>\frac{n[(n-1)^{2}+n]}{(n+1)^{2}}\eqqcolon{\phi}^{*}.$
Any $\phi>\phi^{*}>n-1$ guarantees invertibility of $\mathbf{A}(\mathbf{G})$ and the fact we can express equilibrium efforts as $\mathbf{e}^{*}(\mathbf{G})=(\alpha-\bar{c})\cdot\mathbf{A}(\mathbf{G})^{-1}\mathbf{1}_{n}$. We let $\mathbf{D}$ and $\mathbf{E}$ collect the diagonal and off-diagonal elements of $\mathbf{A}(\mathbf{G})$, respectively, such that for all $i,j\in\mathcal{N}$ with $i\neq j$ we have $D_{ii}=A_{ii}(\mathbf{G})$, $E_{ii}=0$ and $E_{ij}=A_{ij}(\mathbf{G})$. We can then write $\mathbf{A}(\mathbf{G})=\mathbf{D}+\mathbf{E}=(\mathbf{I}_{n}+\mathbf{E}\mathbf{D}^{-1})\mathbf{D}$ and
$\mathbf{e}^{*}(\mathbf{G})=(\alpha-\bar{c})\cdot\mathbf{D}^{-1}(\mathbf{I}_{n}+\mathbf{E}\mathbf{D}^{-1})^{-1}\mathbf{1}_{n}.$
Note that invertibility of $\mathbf{D}$ follows as it is diagonal with $d^{*}>0$ and that of $\mathbf{I}_{n}+\mathbf{E}\mathbf{D}^{-1}$ is guaranteed if $\|\mathbf{E}\mathbf{D}^{-1}\|<1$ for any matrix norm $\|\cdot\|$ (see, e.g., Corollary 5.6.16 of Horn and Johnson, 2012). We will show this holds using the infinity norm defined as $\|\mathbf{A}\|_{\infty}\coloneqq\max_{1\leq i\leq n}\sum_{j=1}^{m}|A_{ij}|$ for any $n\times m$ matrix $\mathbf{A}$. As $\alpha>\bar{c}$, equilibrium efforts $\mathbf{e}^{*}(\mathbf{G})>0$ follows if row sums $\mathbf{s}\coloneqq(\mathbf{I}_{n}+\mathbf{E}\mathbf{D}^{-1})^{-1}\mathbf{1}_{n}$ are entry-wise strictly positive.
First, observe $\|\mathbf{D}^{-1}\|_{\infty}=\max_{i\in\mathcal{N}}|D_{ii}^{-1}|=1/d^{*}$ and $\|E\|_{\infty}=s^{*}$. Under the uniqueness bound $\phi>\phi^{*}$, $\|\mathbf{E}\mathbf{D}^{-1}\|_{\infty}\leq\|\mathbf{E}\|_{\infty}\|\mathbf{D}^{-1}\|_{\infty}=s^{*}/d^{*}<(n-1)^{2}/(n-1)^{2}=1$, proving the invertibility of $\mathbf{I}_{n}+\mathbf{E}\mathbf{D}^{-1}$ and the fact we can write
$(\mathbf{I}_{n}+\mathbf{E}\mathbf{D}^{-1})^{-1}=\sum_{k=0}^{\infty}(-1)^{k}(\mathbf{E}\mathbf{D}^{-1})^{k}=\mathbf{I}_{n}+\sum_{k=1}^{\infty}(-1)^{k}(\mathbf{E}\mathbf{D}^{-1})^{k}.$
Second, using this power series expansion, we can write $\mathbf{s}=\mathbf{1}_{n}+\bm{\varepsilon}$, where we define the remainder term $\bm{\varepsilon}\coloneqq\sum_{k=1}^{\infty}(-1)^{k}(\mathbf{E}\mathbf{D}^{-1})^{k}\mathbf{1}_{n}$. Showing $\|\bm{\varepsilon}\|_{\infty}=\max_{i\in\mathcal{N}}|\varepsilon_{i}|<1$ is then enough to guarantee $s_{i}=1+\varepsilon_{i}>0$ for all $i\in\mathcal{N}$. Observe that under the bound $\phi>\phi^{*}$ we have
$\|\bm{\varepsilon}\|_{\infty}$ $\leq\sum_{k=1}^{\infty}\|(\mathbf{E}\mathbf{D}^{-1})^{k}\|_{\infty}\cdot\|\mathbf{1}_{n}\|_{\infty}\leq\sum_{k=1}^{\infty}\|\mathbf{E}\mathbf{D}^{-1}\|^{k}_{\infty}\leq\sum_{k=1}^{\infty}\left(\frac{s^{*}}{d^{*}}\right)^{k}=\frac{s^{*}}{d^{*}-s^{*}},$
$\implies\|\bm{\varepsilon}\|_{\infty}\leq\frac{n(n-1)^{2}}{(n+1)^{2}\phi-n[n+(n-1)^{2}]}.$
Finally, note that any $\phi$ larger than the threshold $\underline{\phi}$ defined in (8) ensures $\|\bm{\varepsilon}\|_{\infty}<1$. ∎
@@
${e}^{*}_{i}(\mathbf{G}^{C})$ $=\frac{\theta_{i}}{(n+1)^{2}\phi-\theta_{i}^{2}}\left[\alpha-\bar{c}+\sum_{j\neq i}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})\right]=\frac{\theta_{i}}{(n+1)^{2}\phi}\left[\alpha-\bar{c}+\sum_{j=1}^{n}\theta_{j}{e}^{*}_{j}(\mathbf{G}^{C})\right],$
@@ Appendix B Additional Simulation Results
Note: This figure isolates the effect of changing the productivity composition while holding the network fixed. We consider $n=10$ firms partitioned into two fully connected components (two cliques of size 55), mirroring the connectivity pattern of the Positive Assortative (PA) network at $\rho=\tfrac{1}{2}$, and we set $\phi=\underline{\phi}$. Starting from $\rho=0$ (all firms have low productivity $\theta$), we increase $\rho$ in increments of $0.1$ by upgrading one firm at a time from $\theta$ to 11 (high productivity). At each step, we record the change in profit for the transitioning firm (the firm whose productivity is upgraded in that step), holding all other firms’ productivities and the network links fixed. The x-axis reports $\rho$ (the share of high-productivity firms), and the y-axis reports the corresponding profit change of the transitioning firm. Curves are shown for three values of the low productivity parameter, $\theta\in\{0.1,0.5,0.9\}$. Two patterns emerge. First, the profit gain from upgrading is larger when $\theta$ is smaller, consistent with a larger productivity jump from $\theta$ to 11. Second, the profit gain is generally increasing in $\rho$, except around $\rho=\tfrac{1}{2}$, where the fixed network aligns with exact type clustering (all high types concentrated within one clique and all low types within the other); upgrading a low-type firm within the low-type clique yields a smaller marginal benefit than upgrading the remaining low-type firm in the high-type clique. In all cases, the profit change remains positive. Finally, note that this fixed-network exercise is not, in general, a stability analysis: the imposed two-clique network need not be pairwise stable for the corresponding productivity composition, except at $\rho=\tfrac{1}{2}$ (and sufficiently low $\theta$).
@@
Note: Each row corresponds to a different number of firms, ranging from $n=5$ (top row) to $n=200$ (bottom row). Each column corresponds to a different share of high-productivity firms, $\rho$, increasing from $10\%$ (left) to $90\%$ (right). Orange regions indicate parameter combinations for which the PA network is stable, while blue regions indicate stability of the FC network. In light of the sufficient condition for the existence of an equilibrium effort profile (Proposition 1), the cost coefficient $\phi$ on the horizontal axis is normalized by the number of firms. Specifically, since the lower bound satisfies $\underline{\phi}=\mathcal{O}(n)$, the horizontal axis is expressed in terms of $\phi/n$. As a result, within each row the horizontal axis spans the same range, while comparisons across rows account for differences in $n$ through this normalization.